find the first sixth and twelfth terms of the sequence
A(n)=-5+(n-1)(3)

Answer by stanbon(75887)

find the first sixth and twelfth terms of the sequence
A(n)=-5+(n-1)(3)
---
a(1) = -5
----
a(6) = -5+5*3 = 10
----
a(12) = -5+10*3 = 25
======================
Cheers,
Stan H.